A car with the initial velocity of 30 m/s stops at constant rate during a 6 s time interval. the distance traveled by the car during 6 s is:


 


Answer: 90 m


how???????

Grade:12

1 Answers

SHANTANU SHUKLA
12 years ago

U=30M/SEC

V=0

T=6SEC

V=U-AT [RETARDATION]

0=30-A*6

A=30/6

A=5

S=UT+1/2AT2

S=O+1/2*5*36

=90